# Local Topological Constraints on Berry Curvature in Spin–Orbit Coupled BECs

### v2

@@ 4.6 Application to the SOC BEC model
    $\mathcal{P}_{1}(M_{2},g)=\{0\}$, due to the non-triviality of the synthetic gauge potentials $A^{(\pm)}$.
@@
The identification of the BEC metric with the Kaluza–Klein framework follows directly from the construction in Section 2. The first item is a direct consequence of Lemma 4.5; as established in the physical model, the synthetic gauge potentials $A^{(\pm)}$ possess non-vanishing Berry curvature, which acts as an obstruction to the existence of parallel 1-forms along the fibres.
@@ 5 Topological Lower Bound and Its Physical Consequences
With the geometric construction of Section 2 and the cohomological data of Section 3 at hand, we can now apply the general theorem of [12] to the physical connection $\nabla^{C}$. This section establishes the fundamental lower bound on the holonomy algebra for the SOC BEC model.
@@ 5.1 Statement of the bound
The general PT theorem establishes that for a metric connection $\nabla^{C}$ with totally skew-symmetric torsion $T$ on a product Riemannian manifold $M=M_{1}\times M_{2}$, the dimension of the off-diagonal holonomy is constrained by the topology of the torsion class and the geometry of the metric. Specifically, if the condition $\nabla^{LC}T\neq 0$ holds, we have:
@@ 5.2 Main theorem: topological lower bound for SOC BECs
Let $M=T^{2}_{\mathrm{BZ}}\times S^{1}_{\phi_{+}}\times S^{1}_{\phi_{-}}$ be the parameter space equipped with the Kaluza–Klein metric $g_{M}$ and the metric connection $\nabla^{C}$ with torsion $T^{b}=F^{(+)}\wedge A^{(+)}+F^{(-)}\wedge A^{(-)}$. Assuming the Berry curvatures satisfy $F^{(\pm)}]\neq 0$ in $H^{2}(T^{2}_{\mathrm{BZ}};\mathbb{R})$, the following hold:
@@
The proof proceeds by synthesizing the topological and geometric invariants derived in the preceding sections.
1. Mixed tensor rank. As established in Lemma 4.9, the rank $r$ is a purely topological invariant of the class $\omega]=[T^{b}]$. Under the Künneth decomposition for $H^{3}(M;\mathbb{R})$, the torsion class is expressed as $\omega]=[F^{(+)}]\otimes[d\phi_{+}]+[F^{(-)}]\otimes[d\phi_{-}]$. Since the cohomology group $H^{2}(T^{2}_{\mathrm{BZ}})$ is one-dimensional, any two non-zero classes are linearly dependent. Thus, we can write $F^{(-)}]=\lambda[F^{(+)}]$ for some $\lambda\in\mathbb{R}$, which allows the class to be factorized as the simple tensor:
@@
2. Vanishing of the obstruction kernel. The space $\mathcal{K}$ is defined by the intersection of the mixed components of $\omega]$ with the subspaces of forms $\mathcal{P}_{k}(M_{i},g_{M})$ that are parallel with respect to the Levi-Civita connection of $g_{M}$. Crucially, Lemma 4.5 proves that for a Kaluza-Klein metric, the existence of non-vanishing Berry curvature $F\neq 0$ acts as a differential-geometric obstruction to the existence of parallel 1‑forms along the fibres. Consequently, $\mathcal{P}_{1}(S^{1}_{\phi_{\pm}},g_{M})=\{0\}$. Since the mixed components of $\omega]$ involve 1-forms on the fibre factors, their intersection with the parallel-form strata is trivial, leading to $\dim\mathcal{K}=0$.
3. Application of the PT bound. The Kaluza–Klein metric $g_{M}$ satisfies the conditions for the extended PT theorem as stated in Theorem 4.10. Since the non-vanishing Berry curvature acts as a topological obstruction that prevents the existence of global parallel sections, the dimension of the obstruction space $\mathcal{K}$ is zero; consequently, the bound holds directly for the physical Kaluza–Klein metric $g_{M}$ without recourse to auxiliary product metrics. Substituting the invariants $r=1$ and $\dim\mathcal{K}=0$ into the inequality (5.1), we obtain $\dim\mathfrak{hol}^{\mathrm{off}}(\nabla^{C})\geq 1$.
@@ 5.4 Persistence of the bound when the total Chern number vanishes
Let $c_{1}^{(+)}=\frac{1}{2\pi}\int_{T^{2}}F^{(+)}$ and $c_{1}^{(-)}=\frac{1}{2\pi}\int_{T^{2}}F^{(-)}$. If $c_{1}^{(+)}\neq 0$ and $c_{1}^{(-)}=-c_{1}^{(+)}$, then $c_{1}^{\,\mathrm{tot}}=0$. Nevertheless, provided $F^{(+)}]\neq 0$ (hence also $F^{(-)}]\neq 0$), the mixed tensor rank remains $r=1$ and Theorem 5.1 still gives
@@
If $c_{1}^{(+)}\neq 0$ then $F^{(+)}]\neq 0$ in $H^{2}(T^{2})$. The condition $c_{1}^{(-)}=-c_{1}^{(+)}$ implies $F^{(-)}]=-[F^{(+)}]$ in cohomology (because integration gives an isomorphism $H^{2}(T^{2})\cong\mathbb{R}$). Hence both classes are non-zero and the mixed tensor rank attains its maximum $r=1$. The vanishing of $\mathcal{K}$ does not depend on the values of the Chern numbers but only on the non-existence of parallel one-forms on the fibres, which holds as long as $F^{(\pm)}\neq 0$. ∎
This demonstrates that $r^{\sharp}$ detects a local topological constraint that survives when the familiar global invariant vanishes. While the Chern number quantizes the net flux, it can be zero due to cancellation of local contributions; the mixed-rank bound, in contrast, certifies that the curvature cannot be simultaneously flattened along the two independent phase directions.
@@ 5.5 Invariance under metric deformations
The lower bound (5.1) depends only on the topological data $r$ and on the spaces of parallel forms $\mathcal{P}_{k}(M_{i})$. In our construction, these spaces are determined by the Kaluza–Klein metric $g_{M}$, which in turn is built from the physical Berry connections $A^{(\pm)}$. Any smooth deformation of the system that does not change the cohomology classes of $A^{(\pm)}$ and that preserves the vanishing of $\mathcal{P}_{1}(S^{1}_{\phi_{\pm}})$ (i.e., that does not make the Berry curvatures exactly flat) leaves the bound unchanged. This stability makes the result robust against typical experimental perturbations.
@@ 6.2 Locality and robustness beyond the Chern-number paradigm
In configurations where the individual fluxes satisfy $F^{(+)}]\approx-[F^{(-)}]$, the global invariant $c_{1}^{\mathrm{tot}}$ vanishes, potentially suggesting a topologically trivial regime. However, the inequality $\dim\mathfrak{hol}^{\mathrm{off}}\geq 1$ persists as long as the local field strength remains non-zero. This identifies $r^{\sharp}$ as a robust local invariant: it detects the persistence of topological effects, such as the non-integrability of the Berry connection, even in compensated systems or “synthetic vacuum” configurations where net topological charges are absent.
@@ 7.1 Example 1: Rashba–Dresselhaus texture and the maximal rank $r=1$
Consequently, the mixed tensor rank is $r=1$. Given that the obstruction kernel $\mathcal{K}$ vanishes in this geometry, we obtain $r^{\sharp}=1$. Theorem 5.1 then yields:
@@
We note that for a global Rashba texture with non-zero Chern number, the bundle is topologically non-trivial, preventing a global product decomposition. However, applying the theorem within a local trivialization demonstrates that the mixed rank $r^{\sharp}$ correctly identifies the ’locking’ of momentum and phase degrees of freedom. In this context, while the non-vanishing Chern number already implies the existence of vortices on the Brillouin zone, the invariant $r^{\sharp}$ provides an algebraically refined constraint, certifying that the momentum and phase degrees of freedom remain fundamentally coupled through the off-diagonal curvature.
@@ 7.2 Example 2: Topological persistence under vanishing net Chern flux
To illustrate how the bound materializes geometrically, consider the concrete choice
@@
$T=F^{(+)}\wedge A^{(+)}+F^{(-)}\wedge A^{(-)}=\frac{1}{8\pi^{2}}(dk_{x}\wedge dk_{y})\wedge(k_{x}dk_{y}-k_{y}dk_{x}).$
@@ 8 Beyond the Chern Number: Local Rail-Tracks and Prospective Protocols
In the two-component SOC BEC manifold $M=T^{2}_{\mathrm{BZ}}\times S^{1}_{\phi_{+}}\times S^{1}_{\phi_{-}}$, the bound $r^{\sharp}=1$ is non-degenerate since the non-vanishing Berry curvature breaks the integrability of the horizontal distribution, ensuring $\mathcal{K}=0$. Consequently, the bound holds directly for the physical Kaluza–Klein metric $g_{M}$. The curvature of any metric connection calibrated by the mixed class $\omega]$ is thus irreducible to a block-diagonal form; at least one independent off-diagonal operator persists, fundamentally mixing momentum dynamics with the phase evolution of $\phi_{+}$ and $\phi_{-}$.
@@ 9 Conclusions
The extended parameter space $M=T^{2}_{\mathrm{BZ}}\times S^{1}_{\phi_{+}}\times S^{1}_{\phi_{-}}$, equipped with the natural Kaluza–Klein metric $g_{M}$ induced by synthetic gauge fields, carries a metric connection $\nabla^{C}$ with totally skew-symmetric torsion. The harmonic part of this torsion represents a mixed cohomology class $\omega]$ with mixed tensor rank $r=1$ for any generic configuration where the individual Berry curvature classes are non-trivial. Due to the one-dimensionality of $H^{2}(T^{2}_{\mathrm{BZ}})$, these classes are necessarily proportional, forcing $\omega]$ to take the form of a single simple tensor. Crucially, we have shown that the non-vanishing Berry curvature acts as a geometric obstruction to the integrability of the horizontal distribution, ensuring the vanishing of the obstruction kernel $\mathcal{K}=0$. Consequently, the bound $r^{\sharp}=1$ holds directly for the physical metric $g_{M}$.

### v3

@@ 4.5 Application to the SOC BEC model
    The obstruction kernel vanishes: $\mathcal{K}=0$, as shown by Lemma 2.5 under Assumption 2.1.
@@
The identification of the BEC metric with the Kaluza–Klein framework follows directly from the construction in Section 2. The first item follows directly from Lemma 2.5; under Assumption 2.1 the non‑vanishing Berry curvature forces $\mathcal{K}=0$.
@@ 5 Topological Lower Bound and Its Physical Consequences
With the geometric construction of Section 2 and the cohomological data of Section 3 at hand, we can now apply the extended Pigazzini-Toda theorem (Theorem 4.8) to the physical connection $\nabla^{C}$. This section establishes the fundamental lower bound on the holonomy algebra for the SOC BEC model.
@@ 5.1 Statement of the bound
The extended PT theorem (Theorem 4.8) establishes that for a metric connection $\nabla^{C}$ with totally skew-symmetric torsion $T$ on the Kaluza-Klein manifold $M=T^{2}_{\mathrm{BZ}}\times S^{1}_{\phi_{+}}\times S^{1}_{\phi_{-}}$, the dimension of the off-diagonal holonomy is constrained by the topology of the torsion class and the geometry of the metric. Specifically, under Assumption 2.1 and with $\nabla^{LC}T\neq 0$, we have:
@@ 5.2 Main theorem: topological lower bound for SOC BECs
Let $M=T_{\mathrm{BZ}}^{2}\times S_{\phi_{+}}^{1}\times S_{\phi_{-}}^{1}$ be the parameter space equipped with the Kaluza-Klein metric $g_{M}$ (2.1) and the metric connection $\nabla^{C}$ with torsion $T^{b}=F^{(+)}\wedge\Theta^{(+)}+F^{(-)}\wedge\Theta^{(-)}$. Under Assumption 2.1 (constant Berry curvatures), the following hold:
@@
The proof proceeds by synthesizing the topological and geometric invariants derived in the preceding sections and applying the deformation argument developed in Section 4.
1. Mixed tensor rank. As established in Lemma 4.7, the rank $r$ is a purely topological invariant of the class $\omega]=[T^{b}]$. Under the Künneth decomposition for $H^{3}(M;\mathbb{R})$, the torsion class is expressed as $\omega]=[F^{(+)}]\otimes[d\phi_{+}]+[F^{(-)}]\otimes[d\phi_{-}]$. Since the cohomology group $H^{2}(T^{2}_{\mathrm{BZ}})$ is one-dimensional and both $F^{(\pm)}]$ are non-zero under Assumption 2.1, we can write $F^{(-)}]=\lambda[F^{(+)}]$ for some $\lambda=c^{(-)}/c^{(+)}\in\mathbb{R}$, which allows the class to be factorized as the simple tensor:
@@
2. Vanishing of the obstruction kernel. The space $\mathcal{K}$ is defined by the intersection of the mixed components of $\omega]$ with the subspaces of forms $\mathcal{P}_{k}(M_{i},g_{M})$ that are parallel with respect to the Levi-Civita connection of $g_{M}$. Under Assumption 2.1, Lemma 2.5 proves that the obstruction kernel vanishes: $\mathcal{K}=\{0\}$.
3. Application via deformation and lower semi-continuity. Consider the deformation family $\{g_{\varepsilon}\}_{\varepsilon\in[0,1]}$ from Definition 4.1, which interpolates between the product metric $g_{0}$ and the physical Kaluza-Klein metric $g_{M}=g_{1}$. For each $\varepsilon>0$, the metric $g_{\varepsilon}$ satisfies Assumption 2.1 with curvatures scaled by $\varepsilon$. By Lemma 4.2, we have $\mathcal{K}_{\varepsilon}=0$ and $r^{\sharp}_{\varepsilon}=1$ for all $\varepsilon>0$.
Applying Theorem 4.8 (the extension of the PT theorem to Kaluza-Klein metrics with flat fibres) to each $g_{\varepsilon}$ with $\varepsilon>0$, we obtain
$\dim\mathfrak{hol}^{\mathrm{off}}(\nabla^{C_{\varepsilon}})\geq 1\quad\forall\varepsilon>0,$
where $\nabla^{C_{\varepsilon}}$ is the connection built from $g_{\varepsilon}$ with the same torsion form $T$.
By Lemma 4.3, the function $\varepsilon\mapsto\dim\mathfrak{hol}^{\mathrm{off}}(\nabla^{C_{\varepsilon}})$ is lower semi-continuous on $[0,1]$. Therefore,
$\dim\mathfrak{hol}^{\mathrm{off}}(\nabla^{C_{1}})\geq\liminf_{\varepsilon\to 1}\dim\mathfrak{hol}^{\mathrm{off}}(\nabla^{C_{\varepsilon}})\geq 1.$
Since $\nabla^{C_{1}}=\nabla^{C}$ and $g_{1}=g_{M}$, this establishes the bound $\dim\mathfrak{hol}^{\mathrm{off}}(\nabla^{C})\geq 1$ for the physical metric.
@@ 5.4 Persistence of the bound when the total Chern number vanishes
Let $c_{1}^{(+)}=\frac{1}{2\pi}\int_{T^{2}}F^{(+)}$ and $c_{1}^{(-)}=\frac{1}{2\pi}\int_{T^{2}}F^{(-)}$. If $c_{1}^{(+)}\neq 0$ and $c_{1}^{(-)}=-c_{1}^{(+)}$, then $c_{1}^{\,\mathrm{tot}}=0$. Nevertheless, under Assumption 2.1 with $c^{(\pm)}\neq 0$, the mixed tensor rank remains $r=1$ and Theorem 5.1 still gives
@@
If $c_{1}^{(+)}\neq 0$ then $F^{(+)}]\neq 0$ in $H^{2}(T^{2})$. The condition $c_{1}^{(-)}=-c_{1}^{(+)}$ implies $F^{(-)}]=-[F^{(+)}]$ in cohomology (because integration gives an isomorphism $H^{2}(T^{2})\cong\mathbb{R}$). Hence both classes are non-zero and the mixed tensor rank attains its maximum $r=1$. The vanishing of $\mathcal{K}$ does not depend on the values of the Chern numbers but only on the non-existence of parallel one-forms on the fibres, which holds as long as $F^{(\pm)}\neq 0$ (guaranteed by Assumption 2.1). ∎
This corollary is consistent with the geometric setup of Section 2: the global product structure of $M=T^{2}_{\mathrm{BZ}}\times S^{1}_{\phi_{+}}\times S^{1}_{\phi_{-}}$ already implies $c_{1}^{\mathrm{tot}}=0$. The example $c_{1}^{(+)}=-c_{1}^{(-)}$ is therefore not a special case but a generic realization of the trivial-bundle regime under constant curvatures with $\lambda=-1$.
This demonstrates that $r^{\sharp}$ detects a local topological constraint that survives when the familiar global invariant vanishes. While the Chern number quantizes the net flux, it can be zero due to cancellation of local contributions; the mixed-rank bound, in contrast, certifies that the curvature cannot be simultaneously flattened along the two independent phase directions. The lower bound $\dim\mathfrak{hol}^{\mathrm{off}}(\nabla^{C})\geq 1$ thus captures a strictly local topological obstruction that remains active even when global charges cancel.
@@ 5.5 Invariance under metric deformations
The lower bound (5.2) depends only on the topological data $r$ and on the spaces of parallel forms $\mathcal{P}_{k}(M_{i})$. In our construction, these spaces are determined by the Kaluza–Klein metric $g_{M}$, which in turn is built from the physical Berry connections $A^{(\pm)}$. Any smooth deformation of the system that does not change the cohomology classes of $A^{(\pm)}$ and that preserves the vanishing of $\mathcal{P}_{1}(S^{1}_{\phi_{\pm}})$ (i.e., that does not make the Berry curvatures exactly flat) leaves the bound unchanged. This stability is ensured by Lemma 4.5, which guarantees that the non-parallelism condition $\nabla^{LC}T\neq 0$ is open in the space of metrics. This makes the result robust against typical experimental perturbations.
@@ 6.2 Locality and robustness beyond the Chern-number paradigm
In the constant-curvature setting of Assumption 2.1, the local field strength is captured by the constants $c^{(\pm)}$. In configurations where the individual fluxes satisfy $F^{(+)}]\approx-[F^{(-)}]$, the global invariant $c_{1}^{\mathrm{tot}}$ vanishes, potentially suggesting a topologically trivial regime. However, the inequality $\dim\mathfrak{hol}^{\mathrm{off}}\geq 1$ persists as long as both constants $c^{(+)}$ and $c^{(-)}$ are non-zero, regardless of their sum. This identifies $r^{\sharp}$ as a robust local invariant: it detects the persistence of topological effects, such as the non-integrability of the Berry connection, even in compensated systems or “synthetic vacuum” configurations where net topological charges are absent.
@@ 7.1 Example 1: Rashba–Dresselhaus texture and the maximal rank $r=1$
Consequently, the mixed tensor rank is $r=1$. Given that the obstruction kernel $\mathcal{K}$ vanishes in this geometry (as shown in Lemma 2.5 with $\lambda=1$), we obtain $r^{\sharp}=1$. Theorem 5.1 then yields:
@@
Remark 7.1.
For a global Rashba texture with non-zero Chern number, the underlying principal bundle is topologically non-trivial, preventing a global product decomposition. However, as noted in Section 2, our analysis can be applied within local trivializations. The mixed rank $r^{\sharp}$ correctly identifies the ’locking’ of momentum and phase degrees of freedom even in such local patches. While the non-vanishing Chern number already implies the existence of vortices on the Brillouin zone, the invariant $r^{\sharp}$ provides an algebraically refined constraint, certifying that the momentum and phase degrees of freedom remain fundamentally coupled through the off-diagonal curvature.
@@ 7.2 Example 2: Topological persistence under vanishing net Chern flux
To illustrate how the bound materializes geometrically, consider the concrete choice (satisfying Assumption 2.1)
@@
$T=F^{(+)}\wedge d\phi_{+}+F^{(-)}\wedge d\phi_{-}=\frac{1}{2\pi}\,dk_{x}\wedge dk_{y}\wedge(d\phi_{+}-d\phi_{-}).$
@@
The fact that $\dim\mathfrak{hol}^{\mathrm{off}}=2$ in this specific example shows that the bound $\geq 1$ is sharp in the sense that it cannot be improved without additional assumptions, yet particular geometries may yield a larger off-diagonal holonomy algebra.
@@ 8 Beyond the Chern Number: Local Rail-Tracks and Prospective Protocols
In the two-component SOC BEC manifold $M=T^{2}_{\mathrm{BZ}}\times S^{1}_{\phi_{+}}\times S^{1}_{\phi_{-}}$, the bound $r^{\sharp}=1$ is non-degenerate. Under Assumption 2.1, the constant, non-zero Berry curvatures break the integrability of the horizontal distribution, ensuring $\mathcal{K}=0$. Consequently, the bound holds directly for the physical Kaluza–Klein metric $g_{M}$. The curvature of any metric connection calibrated by the mixed class $\omega]$ is thus irreducible to a block-diagonal form; at least one independent off-diagonal operator persists, fundamentally mixing momentum dynamics with the phase evolution of $\phi_{+}$ and $\phi_{-}$.
@@ 9 Conclusions
The extended parameter space $M=T^{2}_{\mathrm{BZ}}\times S^{1}_{\phi_{+}}\times S^{1}_{\phi_{-}}$, equipped with the natural Kaluza–Klein metric $g_{M}$ induced by synthetic gauge fields, carries a metric connection $\nabla^{C}$ with totally skew-symmetric torsion. Under Assumption 2.1 (constant Berry curvatures), the harmonic part of this torsion takes the explicit form given in Lemma 3.1 and represents a mixed cohomology class $\omega]$ with mixed tensor rank $r=1$ for any generic configuration where the individual Berry curvature classes are non-trivial. Due to the one-dimensionality of $H^{2}(T^{2}_{\mathrm{BZ}})$, these classes are necessarily proportional, forcing $\omega]$ to take the form of a single simple tensor. Crucially, we have shown that the non-vanishing Berry curvature acts as a geometric obstruction to the integrability of the horizontal distribution, ensuring the vanishing of the obstruction kernel $\mathcal{K}=0$. Consequently, the bound $r^{\sharp}=1$ holds directly for the physical metric $g_{M}$.
@@ Acknowledgements
Appendix
Appendix A Parallel Vertical 1-Forms
This appendix provides the detailed proof of the condition for vertical 1-forms to be parallel, which is used in the proof of Lemma 2.5. We consider the Kaluza-Klein metric
$g_{M}=\pi^{*}g_{\mathrm{BZ}}+\Theta^{(+)}\otimes\Theta^{(+)}+\Theta^{(-)}\otimes\Theta^{(-)},$
with $\Theta^{(\pm)}=d\phi_{\pm}+\pi^{*}A^{(\pm)}$ and $F^{(\pm)}=dA^{(\pm)}=c^{(\pm)}\operatorname{vol}_{\mathrm{BZ}}$ (Assumption 2.1). Let $\lambda=c^{(-)}/c^{(+)}$.
Orthonormal Frame
Choose coordinates $(x,y)$ on $T^{2}_{\mathrm{BZ}}$ such that $g_{\mathrm{BZ}}=dx^{2}+dy^{2}$ and $\operatorname{vol}_{\mathrm{BZ}}=dx\wedge dy$. An orthonormal frame for $g_{M}$ is given by
$e_{1}=\partial_{x},\quad e_{2}=\partial_{y},\quad e_{3}=\partial_{\phi_{+}},\quad e_{4}=\partial_{\phi_{-}},$
with the corresponding coframe
$e^{1}=dx,\quad e^{2}=dy,\quad e^{3}=\Theta^{(+)},\quad e^{4}=\Theta^{(-)}.$
Connection 1-Forms
The connection 1-forms $\omega^{a}_{\;b}$ for the Levi-Civita connection are defined by
$\nabla^{LC}e_{a}=\omega_{a}^{\;b}\otimes e_{b},$
where $\omega_{ab}=g_{M}(\nabla^{LC}e_{a},e_{b})$ are skew-symmetric. For our Kaluza-Klein metric with flat fibers and constant curvatures, the non-zero connection 1-forms are:
$\omega_{13}$ $=-\omega_{31}=\frac{1}{2}c^{(+)}e^{2},$
$\omega_{14}$ $=-\omega_{41}=\frac{1}{2}c^{(-)}e^{2},$
$\omega_{23}$ $=-\omega_{32}=-\frac{1}{2}c^{(+)}e^{1},$
$\omega_{24}$ $=-\omega_{42}=-\frac{1}{2}c^{(-)}e^{1}.$
All other $\omega_{ab}=0$. This can be derived from the structure equations $de^{a}+\omega^{a}_{\;b}\wedge e^{b}=0$ and the known expressions for $de^{3}$ and $de^{4}$.
Parallel Vertical 1-Forms
A vertical 1-form can be written as
$\eta=h_{+}\Theta^{(+)}+h_{-}\Theta^{(-)}=h_{+}e^{3}+h_{-}e^{4},$
where $h_{\pm}\in C^{\infty}(M)$. The condition for $\eta$ to be parallel, $\nabla^{LC}\eta=0$, becomes
$\nabla^{LC}\eta=(dh_{+}\otimes e^{3}+h_{+}\nabla^{LC}e^{3})+(dh_{-}\otimes e^{4}+h_{-}\nabla^{LC}e^{4})=0.$
Using $\nabla^{LC}e^{3}=-\omega_{3}^{\;b}\otimes e_{b}=\omega_{13}e^{1}\otimes e_{1}+\omega_{23}e^{2}\otimes e_{1}+\omega_{31}e^{1}\otimes e_{2}+\omega_{32}e^{2}\otimes e_{2}$ (and similarly for $e^{4}$), we compute the components of $\nabla^{LC}\eta$.
The components along $e^{1}\otimes e^{3}$ and $e^{2}\otimes e^{3}$ give:
$\frac{\partial h_{+}}{\partial x}=0,\quad\frac{\partial h_{+}}{\partial y}=0,$
so $h_{+}$ is constant along the base. Similarly, from the components along $e^{1}\otimes e^{4}$ and $e^{2}\otimes e^{4}$, we get that $h_{-}$ is also constant along the base.
Now, consider the components along $e^{1}\otimes e^{1}$ and $e^{2}\otimes e^{2}$. These yield:
$h_{+}\omega_{13}(e_{1})+h_{-}\omega_{14}(e_{1})=0,\quad h_{+}\omega_{23}(e_{2})+h_{-}\omega_{24}(e_{2})=0.$
Using the expressions for $\omega_{ab}$, we obtain:
$h_{+}\cdot\frac{1}{2}c^{(+)}+h_{-}\cdot\frac{1}{2}c^{(-)}=0,$
which simplifies to
$h_{+}c^{(+)}+h_{-}c^{(-)}=0\quad\text{or}\quad h_{+}+\lambda h_{-}=0.$
Finally, the vertical derivatives (components along $e^{3}\otimes e^{3}$, etc.) imply that $h_{\pm}$ are also constant along the fibers. Hence, $h_{+}$ and $h_{-}$ are global constants satisfying (A.1).
Therefore, the space of vertical parallel 1-forms is at most one-dimensional, spanned by $\lambda\Theta^{(+)}+\Theta^{(-)}$. However, when expressed in terms of the basis $\{d\phi_{+},d\phi_{-}\}$ of $H^{1}(S^{1}_{\phi_{+}}\times S^{1}_{\phi_{-}})$, a vertical parallel 1-form corresponds to a linear combination $c_{+}d\phi_{+}+c_{-}d\phi_{-}$ such that the constants satisfy (A.1). This completes the derivation of condition ()(*) in the proof of Lemma 2.5.
Appendix B Computation of the Obstruction Map $\Psi_{p}$
This appendix provides the detailed computation of the obstruction map $\Psi_{p}$ defined in Section 4. We work under Assumption 2.1 and use the explicit harmonic representative from Lemma 3.1.
Geometric Setup
Recall the harmonic representative of the torsion class:
$\omega_{h}=\operatorname{vol}_{\mathrm{BZ}}\wedge(a_{+}\Theta^{(+)}+a_{-}\Theta^{(-)}),\quad a_{-}=\lambda a_{+},\quad\lambda=c^{(-)}/c^{(+)}.$
The subspace $\mathcal{V}_{2,1}$ is spanned by the simple tensor
$\xi_{0}=\operatorname{vol}_{\mathrm{BZ}}\otimes(d\phi_{+}+\lambda d\phi_{-}).$
Any element of $\mathcal{V}_{2,1}$ can be written as $\xi=t\xi_{0}$ for some $t\in\mathbb{R}$, which corresponds to taking $a_{+}=t$ and $a_{-}=\lambda t$.
Definition of $\Psi_{p}$
At a point $p=(k,\phi_{+},\phi_{-})\in M$, choose horizontal vectors $X,Y\in\mathcal{H}_{p}$ such that $\operatorname{vol}_{\mathrm{BZ}}(X,Y)=1$. Choose a vertical vector $Z\in\mathcal{V}_{p}$ such that $(d\phi_{+}+\lambda d\phi_{-})(Z)=1$. For $\xi=\operatorname{vol}_{\mathrm{BZ}}\otimes(a_{+}d\phi_{+}+a_{-}d\phi_{-})\in\mathcal{V}_{2,1}$, define
$\Psi_{p}(\xi)=\frac{1}{2}(\nabla^{LC}T)_{\mathrm{off}}(X,Y,Z),$
where $(\cdot)_{\mathrm{off}}$ denotes the off-diagonal part (mapping horizontal to vertical or vice versa). As explained in the main text, the map is well-defined and linear.
Key Computation
We need to compute $(\nabla^{LC}T)_{\mathrm{off}}(X,Y,Z)$. Note that the torsion 3-form $T$ is cohomologous to $\omega_{h}$, but $T$ itself is not necessarily harmonic. However, the definition of $\Psi_{p}$ in the original PT theorem uses the harmonic representative to construct the map. Therefore, we compute using $\omega_{h}$ as the representative of the cohomology class.
Consider the covariant derivative of $\omega_{h}$:
$\nabla^{LC}\omega_{h}=\nabla^{LC}\bigl(\operatorname{vol}_{\mathrm{BZ}}\wedge(a_{+}\Theta^{(+)}+a_{-}\Theta^{(-)})\bigr).$
Since $\operatorname{vol}_{\mathrm{BZ}}$ is parallel on the flat base, we have
$\nabla^{LC}\omega_{h}=\operatorname{vol}_{\mathrm{BZ}}\wedge\bigl(a_{+}\nabla^{LC}\Theta^{(+)}+a_{-}\nabla^{LC}\Theta^{(-)}\bigr).$
From the connection 1-forms computed in Appendix A, we have
$\nabla^{LC}\Theta^{(+)}=\nabla^{LC}e^{3}=\frac{1}{2}c^{(+)}(e^{2}\otimes e^{1}-e^{1}\otimes e^{2}),$
and similarly for $\Theta^{(-)}$ with $c^{(+)}$ replaced by $c^{(-)}$. In coordinate-free notation, this can be written as
$\nabla^{LC}\Theta^{(\pm)}=\frac{1}{2}\pi^{*}F^{(\pm)}.$
Thus,
$\nabla^{LC}\omega_{h}=\operatorname{vol}_{\mathrm{BZ}}\wedge\frac{1}{2}\bigl(a_{+}\pi^{*}F^{(+)}+a_{-}\pi^{*}F^{(-)}\bigr).$
Now, evaluate this 3-form on the vectors $X,Y,Z$. Since $X,Y$ are horizontal and $Z$ is vertical, we have
$(\nabla^{LC}\omega_{h})(X,Y,Z)=\frac{1}{2}\bigl(a_{+}F^{(+)}(\pi_{*}X,\pi_{*}Y)+a_{-}F^{(-)}(\pi_{*}X,\pi_{*}Y)\bigr)\cdot\operatorname{vol}_{\mathrm{BZ}}(X,Y)\cdot 1,$
where we used that $\pi^{*}F^{(\pm)}(X,Y)=F^{(\pm)}(\pi_{*}X,\pi_{*}Y)$ and that $Z$ is vertical so the contraction with the vertical part is trivial. By construction, $\operatorname{vol}_{\mathrm{BZ}}(X,Y)=1$. Moreover, since $F^{(\pm)}=c^{(\pm)}\operatorname{vol}_{\mathrm{BZ}}$, we have
$F^{(\pm)}(\pi_{*}X,\pi_{*}Y)=c^{(\pm)}.$
Therefore,
$(\nabla^{LC}\omega_{h})(X,Y,Z)=\frac{1}{2}(a_{+}c^{(+)}+a_{-}c^{(-)}).$
Off-Diagonal Operator
The off-diagonal part of the curvature operator is then given by raising the index of this 3-form. Specifically, there exists a fixed off-diagonal operator $J:T_{p}M\to T_{p}M$ that exchanges horizontal and vertical directions such that
$(\nabla^{LC}T)_{\mathrm{off}}(X,Y,Z)=(a_{+}c^{(+)}+a_{-}c^{(-)})\,J.$
The precise form of $J$ is not needed for the rank computation; it suffices that $J$ is non-zero and off-diagonal.
Since $a_{-}=\lambda a_{+}$ and $c^{(-)}=\lambda c^{(+)}$, we have
$a_{+}c^{(+)}+a_{-}c^{(-)}=a_{+}c^{(+)}(1+\lambda^{2}).$
Thus,
$\Psi_{p}(\xi)=\frac{1}{2}a_{+}c^{(+)}(1+\lambda^{2})J.$
Kernel and Image
The expression above shows that $\Psi_{p}(\xi)=0$ if and only if $a_{+}=0$, which implies $a_{-}=0$. Hence, $\ker\Psi_{p}=\{0\}$. Moreover, the image of $\Psi_{p}$ is one-dimensional, spanned by the off-diagonal operator $J$. This confirms the claim made in Section 4 that $\dim\operatorname{im}\Psi_{p}=1$.